OpenStudy (anonymous):

Simplify the following expression. 2x+18 divided by x^2+11x+8

sammixboo (sammixboo):

Hate these. You still need help, correct?

OpenStudy (anonymous):

Yes please :)

sammixboo (sammixboo):

I believe it's already simplified o-o

sammixboo (sammixboo):

Unless if you are doing this 

OpenStudy (anonymous):

You're second answer was what I was looking for. Thanks so much for your help!

Nnesha (nnesha):

\[\huge\rm \frac{ 2x+8}{ x^2+11x+8 }\] 1st) check the denominator is it right ? if answer is yes then find common factor of (2x+8) what is common in both terms ? you can't factor x^2+11x+8 it isn't factorable

Nnesha (nnesha):

 there supposed to be plus sign bec x times x = x^2 x +x=2x :)